The perimeter of a rectangle is 46 feet. The length is 5 feet more than the width. Find the dimensions
Answer by Aldorozos(172)

L = length
W = width
L = W +5
Perimeter = 46 = 2(L+W) = 2(W+W+5) =2(2W+5)
46 = 2(2w+5) Dividing both sides by 2 we get 23 = 2W+ 5
Therefore 18 = 2W which means W =9
If W = 9 then L = 9+5 = 14